Circular No. 8017

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002hz IN UGC 12044 M. Ganeshalingam and W. Li, University of California at Berkeley,

report the LOTOSS (cf. IAUC 7906) discovery of an apparent supernova on an unfiltered KAIT image taken on Nov. 16.2 UT (mag ∼ 17.8). The new object was confirmed on an earlier image taken on Nov. 12.2 (mag ∼ 18.4). SN 2002hz is located at α = 22^{h}27^{m}49^{s}.54, δ = +38^{o}35′09′′.5 (equinox 2000.0), which is 13′′.9 east and 2′′.6 south of the nucleus of UGC 12044. A KAIT image taken on Nov. 2.2 showed nothing at this position (limiting mag ∼ 19.0).

SUPERNOVA 2002ia Further to IAUC 8006, W. M. Wood-Vasey, G. Aldering, P. Nugent,

and R. Rudas report the discovery of an apparent supernova (mag 19.2) on unfiltered NEAT images taken at Haleakala on Oct. 30.52, 30.53, and 30.54 UT. The new object was confirmed on NEAT images taken at Palomar on Nov. 13 (at mag 18.6) and 14 (mag 18.5). SN 2002ia is located at α = 1^{h}45^{m}38^{s}.52, δ = +24^{o}08′07′′.9 (equinox 2000.0), which is 4′′ west and 4′′ north of the center of the apparent host galaxy. No point source is visible at this location in previous NEAT images taken on 2001 Nov. 11 (limiting mag 20.0 at S/N = 5).

2002 TD_{60}

P. Pravec and L. Šarounová, Ondřejov Observatory; C. Hergenrother, Lunar and Planetary Laboratory; P. Brown and G. Esquerdo, University of Western Ontario; G. Masi, C. Belmonte, and F. Mallia, Campo Catino Observatory; and A. W. Harris, Space Science Institute, La Canada, CA, communicate: “Our lightcurve observations of 2002 TD_{60} (MPEC 2002-T33, 2002-V32) on seven nights during Nov. 7.1–14.1 UT revealed a complex lightcurve that shows two periods of 2.00746 and 2.8507 hr (formal errors 0.00005 and 0.0002 hr, respectively), indicating that the object is in a nonprincipal axis-rotation state. The large amplitude of the lightcurve (2.0 mag) indicates that the body is quite elongated, with the maximum-to-minimum axis ratio ∼ 3. The fast-tumbling, elongated body cannot be held together by self-gravitation only and must be a coherent body with nonzero tensile strength for a plausible asteroidal bulk density of < 8 g/cm^{3}. Further observations, especially with different techniques, are very desirable to describe the unique object.”

2002 November 18  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green